solve for the electric flux that acts on a uniform electric field with a magnitude of 8000N/C approximately passes through a 10m^2 flat surface area


1
Expert's answer
2022-02-20T15:46:43-0500

Answer

the electric flux is given by

"\phi=EA\cos\theta"

Putting all values

"\phi=800*10*\cos0°\\=8000Wb."